OpenStudy (anonymous):

one jug holds 7/8 gallon of water. If a faucet can release 16 gallons of water in 8 minutes,how long would it take to fill one jug with water

OpenStudy (anonymous):

8/16 * 7/8 = 7/16 minutes

OpenStudy (anonymous):

=7/16 minutes